Find the square of the following numbers. 
32
Find the square of the following numbers. 
32

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(32)^2
= (30+2 )^2
= (30)^2+ (2)^2+ 2×30× [Since, (a+b)^2= a^2+b^2+2ab]​
= 900+4+120
= 1024

Thanks